OpenStudy (priyar):

How many real solutions does the equation x^7+14x^5+16x^3+30x-560=0 have?

OpenStudy (roadjester):

What kind of approach are you looking for? A solution by hand, or by calculator?

OpenStudy (priyar):

by hand of course ! and i want to know how to solve this fat...shortcut maybe

OpenStudy (priyar):

*fast

OpenStudy (anonymous):

x^7+14x^5+16x^3+30x-560=0 this is your problem yeah? Okay, first take 560 away on both sides, now you have x^7+14^5+16^3+30x = 560 yeah? Okay, now that you have simplified and solved the problem a little now divide both sides by 30. So you now have x^7+14x^5+16x^3+x = 18.6 repeating yeah? Since both sides share a common factor of 2, then divide both sides by 2. Now you have x^7+7x^5+8x^3+x = 9.3. You can't simplify this anymore.

OpenStudy (priyar):

if we divide both sides by 30 X^7 will become x^7/30 and so on for all the terms u forgot to write that...

OpenStudy (roadjester):

The most significant coefficient is 1, not 30.... @StudiousHojea

OpenStudy (roadjester):

As for the number of real solutions, I just graphed the equation, there are none.

OpenStudy (anonymous):

A polynomial equation of nth degree has n solutions, no?

OpenStudy (priyar):

yes @aleroth but all need not be real right?

OpenStudy (anonymous):

If I recall correctly, it has n real solutions

OpenStudy (anonymous):

So the answer should be 7

OpenStudy (roadjester):

nope, it has 7 possible, but none are real

OpenStudy (priyar):

Actually, that is not correct....for ex:even qudratic eq can have non-real roots...that is why we have...discriminant...remember @aleroth D=b^2-4ac...

OpenStudy (roadjester):

The graph show that it never crosses the x-axis, therefore no real solutions.

OpenStudy (priyar):

are u sure @roadjester ?

OpenStudy (anonymous):

Oops, copied in the equation wrong, but I see 1 real solution

OpenStudy (anonymous):

Are you supposed to use a graph drawer for this?

OpenStudy (priyar):

I want to know how to solve algebraically..not through graph...

OpenStudy (roadjester):

Well, what class is this for? So I know what I can and can't use.

OpenStudy (priyar):

11th

OpenStudy (roadjester):

I mean, is this for Algebra 1, Algebra 2, Pre-Calculus, etc

OpenStudy (anonymous):

Okay, my bad, what I meant was divide both sides by 14 so that the problem will be like this: x^7+14x^5+16x^3+30x = 40 ---------------------------- 14 Do you see where this going? The 14 cancels out the other 14 so you are left with: x^7+x^5+16x^3+30x = 40. Then you divide both sides by 10 so that the 30 becomes 3 and the 40 becomes 4 like this. So this is how the equation will look x^7+x^5+16x^3+3x = 4. Then you divide both sides by 4 so that the 16 in the equation will be reduced to 4 so this is how the equation looks now: x^7+X^5+4x^3+3x = 1. Then divide both sides by x. This is how the equation looks now: x^7+x^5+4x^3+3 = 1/x. I got this because I cancelled out the x in the equation but not the ones with the exponents.

OpenStudy (priyar):

Does anyone know Descarter's rule of signs?

OpenStudy (roadjester):

@StudiousHojea I have no clue what you're trying to do since you're not being consistent in your division of terms. Your x^7 term isn't divided. O.o

OpenStudy (anonymous):

That's the thing, we can't do anything about the variables with the exponents.

OpenStudy (priyar):

ok thanks all...i found the answer

OpenStudy (roadjester):

And as for descartes rule of signs, i remember learning bout it when I was in high school. Is that what you're supposed to use?

OpenStudy (anonymous):

Oh, that's great!

OpenStudy (roadjester):

Ok, sweet! Good luck in your studies!

OpenStudy (priyar):

thank u

OpenStudy (anonymous):

Sorry roadjester, I like math but people tell me I am not very good at it so I was just telling her the stuff I learned in Algebra 1.